# Proximity-induced superconductivity and emerging topological phases in altermagnet-based heterostructures

### v1

@@ IV Pairing Symmetry with RSOC
Figure 5: Panels (a)–(d) demonstrate the frequency dependence of the pairing amplitudes at an arbitrary fixed momentum point $(k_{x},k_{y})=(1,0.4)$ (in units of the inverse lattice constant) in the Brillouin zone. Panel (a) displays the real and imaginary parts of the singlet component, panels (b) and (c) present the in-plane triplet components $d_{x}$, $d_{y}$, and panel (d) illustrates the transverse triplet component $d_{z}$. In panel (e), we exhibit the absolute magnitudes of the triplet components as a function of the Rashba SOC strength $\alpha$ (the inset depicts the singlet component), while panel (f) presents the absolute magnitudes of both singlet and triplet components as a function of the AM-SC coupling strength $\lambda_{s}$, plotted on a logarithmic scale. The other model parameters are chosen as $J_{a}=0.1\Delta$, $\mu=0t$, $\Delta=0.3t$, $\lambda_{s}=0.5\Delta$.
@@
where, $\alpha$ is the Rashba coupling strength. Following the previous section, we obtain the analytic expressions for the singlet and triplet components of the pairing amplitudes as
@@ V Comparison with Exact Diagonalization Results
    When both the inter-layer hopping and the altermagnetic exchange field are absent (i.e.,  $\tilde{t}=0$ and $J_{a}=0$), the AM layer exhibits a metallic DOS with a peak at zero energy akin to 2D tight-binding model (see Fig. 6(a)). This result is analogous to the $\lambda_{s}=0$ case shown in Fig. 2(a).
@@
    Introducing a finite inter-layer hopping amplitude ($\tilde{t}\neq 0$) opens a superconducting gap around zero energy in the AM layer, indicative of proximity-induced superconductivity (see Fig. 6(a)). This mimics the finite $\lambda_{s}$ behavior presented in Fig. 2(a).
@@ VI Emergence of Topological Superconducting Phases
Figure 7: Panels (a)–(c) exhibit the edge spectra employing OBC along the $x$-axis and PBC along the $y$-axis for three different coupling strengths: $\lambda_{s}=0$, $0.2\Delta$, and $0.5\Delta$. Panel (d) represents the site-resolved LDOS distribution at zero energy with OBC along both $x$ and $y$-directions. We consider a finite square lattice with $60\times 60$ lattice sites at $\lambda_{s}=0.5\Delta$. The inset displays a zoomed-in view of the quasi-energy spectrum near the zero-energy. The other model parameters are used as $t=1$, $\mu=-0.4t$, $\Delta=0.3t$, $b=1$, $\alpha=0.3t$, and $J_{a}=0.3t$.
We have already seen that the inclusion of RSOC leads to the generation of odd-parity $p$-wave pairing amplitudes, $d_{x}$ and $d_{y}$ in the ETO class (see Table 1), and therefore we expect the emergence of topological phases in our system. To further investigate about the topological superconducting properties of the proximity induced altermagnetic layer, we consider the following Hamiltonian for our system,
$H(k_{x},k_{y})$ $=[\xi(\textbf{k})+\mu]\,\tau_{z}\sigma_{0}+h_{AM}(\textbf{k})\,\tau_{0}\sigma_{z}$
@@
where, $\xi(\mathbf{k})=-2t(\cos k_{x}+b\cos k_{y})$. We introduce a dimensionless parameter $0<b<1$ to represent the anisotropic hopping distortion along the $y$-direction. The last term $\Sigma(\omega=0)$ in Eq. (22) represents the self-energy correction arising from the proximity-induced SC as discussed before.
@@
To better understand the origin of the MEMs even when the Chern number vanishes, we construct an effective 1D Hamiltonian along the high-symmetry line $k_{y}=0$. For this 1D model, which preserves chiral symmetry $S=\tau_{y}\sigma_{y}$, the topological properties can be characterized by an integer winding number $W$ Ryu et al. (2010); Chiu et al. (2016). In the chiral basis where $S$ is diagonal, the Hamiltonian takes an off-diagonal form,
@@
Figure 8: Panel (a) shows the winding number $W$ in the $J_{a}$$\lambda_{s}$ plane for the symmetric case ($b=1$), calculated at fixed $k_{y}=0$ with $\mu=-0.4t$. Panel (b) depicts the edge spectrum with OBC along the $x$-axis and PBC along the $y$-axis for the anisotropic case ($b=0.85t$), where the gap-closing points at $k_{y}=\pi$ are gapped out. The chosen parameters are $\mu=-0.4t$ and $\lambda_{s}=0.5\Delta$. Panel (c) displays the Chern number for the anisotropic case in the $\mu$$\lambda_{s}$ plane with $J_{a}=0.3t$, while panel (d) exhibits the Chern number in the $J_{a}$$\lambda_{s}$ plane with $\mu=-0.4t$. Unless otherwise specified, the model parameters are $\Delta=0.3t$ and $\alpha=0.3\Delta$.
The coexistence of a zero Chern number ($C=0$) with a finite winding number ($W=1$) indicates that the system supports a weak topological superconducting (WTSC) phase. Using the effective Hamiltonian in Eq. (22), we qualitatively reinvestigate the predicted transition from a weak to a strong topological superconducting phase (STSC) with $\lvert C\rvert=1$, as suggested by the phenomenological model Ghorashi et al. (2024) and Eq. (37) of Appendix C. We confirm that incorporating a small anisotropic hopping term ($b=0.85$) gaps out the helical edge states at $k_{y}=\pm\pi$, while those at $k_{y}=0$ remain gapless dispersive, resulting in a finite Chern number of $\lvert C\rvert=1$. This transition is illustrated in the edge spectra of Fig. 8(b). Moreover, the phase diagrams depicted in Figs. 8(c) and (d) demonstrate the same via the appearance of finite Chern number $\lvert C\rvert=1$ in $\mu$$\lambda_{s}$ and $J_{a}$$\lambda_{s}$ planes respectively.
@@ VII Summary and Conclusions
To summarize, in this article, we investigate superconducting proximity effects in AM-SC heterostructures using a microscopic approach. Our setup consists of a 2D altermagnetic layer placed on the surface of a three-dimensional $s$-wave SC, with direct tunneling between the first layer of the SC and the AM layer. By integrating out the superconducting degrees of freedom, we derive an effective Green’s function and a corresponding low-energy effective Hamiltonian for the AM layer. Using this framework, we analyze the system’s spectral features, demonstrating that a superconducting gap is indeed induced in the AM layer. The magnitude of this proximity induced gap depends on the interface coupling strength and the normal-state density of states of the SC. The resulting DOS exhibits spin-split peaks alongside characteristic superconducting coherence peaks. Furthermore, an analysis of the anomalous part of the Green’s function reveals that the AM induces out-of-plane spin-triplet pairing correlations with even parity. Introducing RSOC into the AM layer generates additional in-plane triplet components with odd parity, enabling the emergence of TSC phase. We further show that this effective Hamiltonian can host both WTSC and STSC phase, characterized by the presence of localized MEMs. The topological nature of these phases is established via the Chern number calculations. We also analyze the agreement of our effective Hamiltonian based results with those obtained from numerical ED.
In literature, there exist numerous studies on the superconducting proximity effect across a wide range of magnetic and non-magnetic systems. On the magnetic side, it has been extensively explored in ferromagnets (FMs) Buzdin (2005); Halterman and Valls (2001); Salamone et al. (2022); Efetov et al. (2008), inhomogeneous FM Linder et al. (2009), SC–FM–SC heterostructures Golovchanskiy et al. (2020), and antiferromagnetic systems Bobkov et al. (2023); Fyhn et al. (2023). The superconducting proximity effect has also been investigated in various non-magnetic modern platforms, such as topological insulators, Rashba nanowires, Weyl semimetals, graphene, and various mesoscopic junctions Burset et al. (2015); Fu and Kane (2008); Khanna et al. (2014); Chevallier et al. (2013); Moriya et al. (2020); Sitthison and Stanescu (2014); Linder et al. (2010); Black-Schaffer and Balatsky (2013a).
@@ Acknowledgments
O.A., A.P., and A.S. acknowledge SAMKHYA: High-Performance Computing Facility provided by Institute of Physics, Bhubaneswar and the two Workstations provided by Institute of Physics, Bhubaneswar from DAE APEX Project, for numerical computations. P.D. acknowledges the Department of Space (DoS), Government of India, for all support at PRL. O.A. acknowledges Kamalesh Bera for stimulating discussions.
@@ Appendix A Derivation of the Self-Energy
$S=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\bigg[$ $\sum_{\mathbf{R,R^{\prime}}}\tilde{\Phi}_{\mathbf{R}}^{\dagger}\,\left[G_{\text{SC}}(\omega)^{-1}_{\mathbf{R,R^{\prime}}}\right]\tilde{\Phi}_{\mathbf{R^{\prime}}}$
@@
$\Sigma_{r,r^{\prime}}(\omega)=\sum_{\mathbf{R,R^{\prime}}}A_{r,\mathbf{R}}\,G_{\text{SC}}(\omega)_{\mathbf{R,R^{\prime}}}\,A_{\mathbf{R^{\prime}},r^{\prime}}^{\dagger}\ .$ (32)
@@
$G_{\text{SC}}(\omega,\mathbf{k})=\frac{1}{\omega^{2}\text{-}\xi_{\mathbf{k}}^{2}\text{-}\Delta^{2}}\begin{pmatrix}\omega\text{+}\xi_{\mathbf{k}}&0&\Delta&0\\ 0&\omega\text{+}\xi_{\mathbf{k}}&0&\Delta\\ \Delta&0&\omega\text{-}\xi_{\mathbf{k}}&0\\ 0&\Delta&0&\omega\text{-}\xi_{\mathbf{k}}\end{pmatrix}\,$ (33)
@@
$G_{\text{SC}}(\omega)_{\mathbf{R,R^{\prime}}}=\frac{1}{N}\sum_{\mathbf{k}}G_{\text{SC}}(\omega,\mathbf{k})\,e^{i\mathbf{k}\cdot(\mathbf{R}-\mathbf{R^{\prime}})}\ .$ (34)
@@
$\Sigma_{r_{c}}(\omega)=|\tilde{t}|^{2}\,G_{\text{SC}}(\omega)_{\mathbf{R}_{c},\mathbf{R}_{c}}$
@@ Appendix C Phenomenological Hamiltonian and Green’s Function
$H(k_{x},k_{y})=-2t(\cos k_{x}+b\cos k_{y})\tau_{0}\sigma_{0}+\mu\tau_{z}\sigma_{0}+J_{a}(\cos k_{x}-\cos k_{y})\tau_{0}\sigma_{z}+2\alpha(\sin k_{y}\tau_{z}\sigma_{x}-\sin k_{x}\tau_{z}\sigma_{y})+\Delta\tau_{x}\sigma_{0}\ ,$ (37)
@@ Appendix D Analysis of the pairing symmetry for $p$-wave magnets
Here, we explore the symmetry of the pair amplitudes proximitized to a $p$-wave magnet to elucidate the odd-parity nature of its triplet components. The model Hamiltonian for the $p$-wave magnet is given by Brekke et al. (2024); Nagae et al. (2025); Hellenes et al. ; Maeda et al. (2024a),

### v2

@@ IV Pairing Symmetry with RSOC
Figure 5: Panels (a)–(d) demonstrate the frequency dependence of the pairing amplitudes at an arbitrary fixed momentum point $(k_{x},k_{y})=(1,0.4)$ (in units of the inverse lattice constant) in the Brillouin zone. Panel (a) displays the real and imaginary parts of the singlet component, panels (b) and (c) present the in-plane triplet components $d_{x}$, $d_{y}$, and panel (d) illustrates the transverse triplet component $d_{z}$. In panel (e), we exhibit the absolute magnitudes of the triplet components as a function of the Rashba SOC strength $\alpha$ (the inset depicts the singlet component), while panel (f) presents the absolute magnitudes of both singlet and triplet components as a function of the AM-SC coupling strength $\lambda_{s}$, plotted on a logarithmic scale. The other model parameters are chosen as $J_{a}=0.1\Delta$, $\mu=0$, $\Delta=0.3t$, $\lambda_{s}=0.5\Delta$.
@@
where, $\alpha$ is the RSOC strength. Note that, the RSOC term in Eq. (16) is incorporated assuming an interfacial layer between the AM and SC, not intrinsic to AM which would compromise the altermagnetic classification. However, for simplicity, we model the combined effect of AM and RSOC by the effective tight-binding Hamiltonian $H_{\mathrm{AM+RSOC}}(\mathbf{k})$ instead of considering a microscopic tunneling between the AM and RSOC layer, or, considering the presence of RSOC in the SC layer itself Heinsdorf and Franz (2026). Following the previous section, we obtain the analytic expressions for the singlet and triplet components of the pairing amplitudes as
@@ V Comparison with Exact Diagonalization Results
    When both the inter-layer hopping and the altermagnetic exchange field are absent (i.e.,  $\tilde{t}=0$ and $J_{a}=0$), the AM layer exhibits a metallic DOS with a peak at zero-energy akin to 2D tight-binding model (see Fig. 6(a)). This result is analogous to the $\lambda_{s}=0$ case shown in Fig. 2(a).
@@
    Introducing a finite inter-layer hopping amplitude ($\tilde{t}\neq 0$) opens a superconducting gap around zero energy in the AM layer, indicative of proximity-induced superconductivity (see Fig. 6(a)). This mimics the finite $\lambda_{s}$ behavior presented in Fig. 2(a). Interestingly, an additional pair of coherence peaks with lower amplitude appears in Fig. 6(a). This corresponds to the coherence peaks of the parent SC and can appear in the AM as higher order process. These additional peaks are absent in Fig. 2(a) where the SC degrees of freedom are integrated out as described using the Green’s-function approach in the tunneling regime (see Appendix A for details).
@@ VI Emergence of Topological Superconducting Phases
Figure 7: Panels (a)–(c) exhibit the edge spectra employing OBC along the $x$-axis and PBC along the $y$-axis for three different coupling strengths: $\lambda_{s}=0$, $0.2\Delta$, and $0.5\Delta$. Panel (d) represents the site-resolved LDOS distribution at zero energy with OBC along both $x$ and $y$-directions. We consider a finite square lattice with $60\times 60$ lattice sites at $\lambda_{s}=0.5\Delta$. The inset displays a zoomed-in view of the quasi-energy spectrum near the zero-energy. The other model parameters are used as $t=1$, $\mu=0.4t$, $\Delta=0.3t$, $b=1$, $\alpha=0.3t$, and $J_{a}=0.3t$.
We have already seen that the inclusion of RSOC leads to the generation of odd-parity $p$-wave pairing amplitudes, $d_{x}$ and $d_{y}$ in the ETO class (see Table 1), and therefore we expect the emergence of topological phases in our system. To further investigate about the topological superconducting properties of the proximity-induced AM layer, we consider the following Hamiltonian for our system,
$H(k_{x},k_{y})$ $=[\xi(\textbf{k})-\mu]\,\tau_{z}\sigma_{0}+h_{AM}(\textbf{k})\,\tau_{0}\sigma_{z}$
@@
where, $\xi(\mathbf{k})=-2t(\cos k_{x}+b\cos k_{y})$. We introduce a dimensionless parameter $0<b<1$ to represent the anisotropic hopping distortion along the $y$-direction. The last term $\Sigma(\omega=0)$ in Eq. (22) represents the self-energy correction arising from the proximity-induced SC as discussed before. Interestingly, the zero-frequency ($\omega=0$) self-energy correction $\Sigma(0)=\frac{\lambda_{s}\Delta\tau_{x}}{\sqrt{\Delta^{2}+\eta^{2}}}$ appears to be independent of $\Delta$ in the limit $\eta=0$. However, this result is not counter-intuitive as ${\rm lim}_{\eta\rightarrow 0}\Sigma(0)=\lambda_{s}\tau_{x}$ depends on $\Delta$ indirectly and can be understood as following. If we start with $\Delta=0$ and then consider the limit $\omega\rightarrow 0$, then $\Sigma(0)$ vanishes identically. Thus, there will be no proximity-induced superconducting gap in the AM layer for $\Delta=0$ which is physically correct. Therefore, ${\rm lim}_{\eta\rightarrow 0}\Sigma(0)=\lambda_{s}\tau_{x}$ is valid only for $\Delta\neq 0$. In addition, the proximity-induced gap is significantly controlled by the coupling strength $\lambda_{s}$. In the expression, ${\rm lim}_{\eta\rightarrow 0}\Sigma(0)=\lambda_{s}\tau_{x}$, $\lambda_{s}$ effectively mimicks the proximity-induced superconducting gap, given a nonzero value of parent gap $\Delta$ Stanescu and Das Sarma (2017).
@@
To better understand the origin of the MEMs even when the Chern number vanishes, we construct an effective 1D Hamiltonian along the high-symmetry line $k_{y}=0$. For this 1D model, which preserves chiral symmetry $S=\tau_{y}\sigma_{y}$, the topological properties can be characterized by an integer winding number $W$ Ryu et al. (2010); Chiu et al. (2016); Pal and Ghosh (2025). In the chiral basis where $S$ is diagonal, the Hamiltonian takes an off-diagonal form,
@@
Figure 8: Panel (a) shows the winding number $W$ in the $J_{a}$$\lambda_{s}$ plane for the symmetric case ($b=1$), calculated at fixed $k_{y}=0$ with $\mu=0.4t$. Panel (b) depicts the edge spectrum with OBC along the $x$-axis and PBC along the $y$-axis for the anisotropic case ($b=0.85t$), where the gap-closing points at $k_{y}=\pi$ are gapped out. The chosen parameters are $\mu=0.4t$ and $\lambda_{s}=0.5\Delta$. Panel (c) displays the Chern number for the anisotropic case in the $\mu$$\lambda_{s}$ plane with $J_{a}=0.3t$, while panel (d) exhibits the Chern number in the $J_{a}$$\lambda_{s}$ plane with $\mu=0.4t$. Unless otherwise specified, the model parameters are $\Delta=0.3t$ and $\alpha=0.3t$.
The coexistence of a zero Chern number ($C=0$) with a finite winding number ($W=1$) indicates that the system supports a weak topological superconducting (WTSC) phase Zhu et al. (2024). Using the effective Hamiltonian in Eq. (22), we qualitatively reinvestigate the predicted transition from a weak to a strong topological superconducting phase (STSC) with $\lvert C\rvert=1$, as suggested by the phenomenological model Ghorashi et al. (2024); Zhu et al. (2024) and Eq. (37) of Appendix C. We confirm that incorporating a small anisotropic hopping term ($b=0.85$) gaps out the helical edge states at $k_{y}=\pm\pi$, while those at $k_{y}=0$ remain gapless dispersive, resulting in a finite Chern number of $\lvert C\rvert=1$. This transition is illustrated in the edge spectra of Fig. 8(b). Moreover, the phase diagrams depicted in Figs. 8(c) and (d) demonstrate the same via the appearance of finite Chern number $\lvert C\rvert=1$ in $\mu$$\lambda_{s}$ and $J_{a}$$\lambda_{s}$ planes respectively.
@@ VII Summary and Conclusions
To summarize, in this article, we investigate superconducting proximity effects in AM-SC heterostructures using a microscopic approach. Our setup consists of a 2D altermagnetic layer placed on the surface of a three-dimensional $s$-wave SC, with direct tunneling between the first layer of the SC and the AM layer. By integrating out the superconducting degrees of freedom, we derive an effective Green’s function and a corresponding low-energy effective Hamiltonian for the AM layer. Using this framework, we analyze the system’s spectral features, demonstrating that a superconducting gap is indeed induced in the AM layer. The magnitude of this proximity induced gap depends on the interface coupling strength and the normal-state density of states of the SC. The resulting DOS exhibits spin-split peaks alongside characteristic superconducting coherence peaks. Furthermore, an analysis of the anomalous part of the Green’s function reveals that the AM induces out-of-plane spin-triplet pairing correlations with even parity. Introducing RSOC into the AM layer generates additional in-plane triplet components with odd parity, enabling the emergence of TSC phase. We further show that this effective Hamiltonian can host both WTSC and STSC phase, characterized by the presence of localized MEMs. The topological nature of these phases is established via the Chern number calculations. We also analyze the agreement of our effective Hamiltonian based results with those obtained from numerical ED. We note that in this work, we have considered two band model for the altermagnetic spin-splitting. Nevertheless, one can also consider a four band model Šmejkal et al. (2022a); Rasmussen et al. (2025) to investigate the multi-orbital effects and their role in enhancing the induced pairing correlations. Here, we restrict ourselves to the two band model and will investigate the orbital effect in future.
In literature, there exist numerous studies on the superconducting proximity effect across a wide range of magnetic and non-magnetic systems. On the magnetic side, it has been extensively explored in ferromagnets (FMs) Buzdin (2005); Halterman and Valls (2001); Salamone et al. (2022); Efetov et al. (2008), inhomogeneous FM Linder et al. (2009), SC–FM–SC heterostructures Golovchanskiy et al. (2020), and antiferromagnetic systems Bobkov et al. (2023); Fyhn et al. (2023). Although, within the present framework AM–SC and FM–SC heterostructures belong to the same symmetry classes as the magnetic order parameter is even under parity for both of them. However, the resulting proximity-induced pairing correlations exhibit qualitatively different momentum-space structures reflecting the underlying magnetic order (see Appendix. E for more detailed comparision). The superconducting proximity effect has also been investigated in various non-magnetic modern platforms, such as topological insulators, Rashba nanowires, Weyl semimetals, graphene, and various mesoscopic junctions Burset et al. (2015); Fu and Kane (2008); Khanna et al. (2014); Chevallier et al. (2013); Moriya et al. (2020); Sitthison and Stanescu (2014); Linder et al. (2010); Black-Schaffer and Balatsky (2013a).
@@ Acknowledgments
We acknowledge A. M. Black-Schaffer for useful comments and discussions. O.A. acknowledges K. Bera for stimulating discussions. O.A., A.P., and A.S. acknowledge SAMKHYA: High-Performance Computing Facility provided by Institute of Physics, Bhubaneswar and the two Workstations provided by Institute of Physics, Bhubaneswar from DAE APEX Project, for numerical computations. The work of PD at Physical Research Laboratory was supported by the Department of Space, Govt. of India.
@@ Appendix A Derivation of the Self-Energy
$S=\int_{-\infty}^{\infty}\frac{d\omega}{2\pi}\bigg[$ $\sum_{\mathbf{R,R^{\prime}}}\tilde{\Phi}_{\mathbf{R}}^{\dagger}\,\left[G^{0}_{\text{SC}}(\omega)^{-1}_{\mathbf{R,R^{\prime}}}\right]\tilde{\Phi}_{\mathbf{R^{\prime}}}$
@@
$\Sigma_{r,r^{\prime}}(\omega)=\sum_{\mathbf{R,R^{\prime}}}A_{r,\mathbf{R}}\,G^{0}_{\text{SC}}(\omega)_{\mathbf{R,R^{\prime}}}\,A_{\mathbf{R^{\prime}},r^{\prime}}^{\dagger}\ .$ (32)
@@
$G^{0}_{\text{SC}}(\omega,\mathbf{k})=\frac{1}{\omega^{2}\text{-}\xi_{\mathbf{k}}^{2}\text{-}\Delta^{2}}\begin{pmatrix}\omega\text{+}\xi_{\mathbf{k}}&0&\Delta&0\\ 0&\omega\text{+}\xi_{\mathbf{k}}&0&\Delta\\ \Delta&0&\omega\text{-}\xi_{\mathbf{k}}&0\\ 0&\Delta&0&\omega\text{-}\xi_{\mathbf{k}}\end{pmatrix}\,$ (33)
@@
$G^{0}_{\text{SC}}(\omega)_{\mathbf{R,R^{\prime}}}=\frac{1}{N}\sum_{\mathbf{k}}G^{0}_{\text{SC}}(\omega,\mathbf{k})\,e^{i\mathbf{k}\cdot(\mathbf{R}-\mathbf{R^{\prime}})}\ .$ (34)
@@
$\Sigma_{r_{c}}(\omega)=|\tilde{t}|^{2}\,G^{0}_{\text{SC}}(\omega)_{\mathbf{R}_{c},\mathbf{R}_{c}}$
@@ Appendix C Phenomenological Hamiltonian and Green’s Function
$H(k_{x},k_{y})=-2t(\cos k_{x}+b\cos k_{y})\tau_{0}\sigma_{0}-\mu\tau_{z}\sigma_{0}+J_{a}(\cos k_{x}-\cos k_{y})\tau_{0}\sigma_{z}+2\alpha(\sin k_{y}\tau_{z}\sigma_{x}-\sin k_{x}\tau_{z}\sigma_{y})+\Delta\tau_{x}\sigma_{0}\ ,$ (37)
@@ Appendix D Analysis of the pairing symmetry for $p$-wave magnets
Here, we explore the symmetry of the pair amplitudes proximitized to a $p$-wave magnet to elucidate the odd-parity nature of its triplet components. The model Hamiltonian for the $p$-wave magnet is given by Brekke et al. (2024); Nagae et al. (2025); Hellenes et al. ; Maeda et al. (2024),